OpenStudy (anonymous):

Hamburgers cost $2.50 and cheeseburgers cost $3.50 at a snack bar. Ben has sold no more than $30 worth of hamburgers and cheeseburgers in the first hour of business. Let x represent the number of hamburgers and y represent the number of cheeseburgers. The inequality 2.50x + 3.50y ≤ 30 represents the food sales in the first hour. If Ben has sold 4 cheeseburgers, what is the maximum value of hamburgers Ben could have sold?

OpenStudy (anonymous):

@Secret-Ninja , @iGreen

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@ganeshie8

OpenStudy (secret-ninja):

Okay, first, you have to change the variable for the cheeseburgers (y) to 4 because he sold 4. You now have: 2.50x + 3.50 x 4 <= 30 So now you just solve for x. \[ 2.50x + 3.50 \times 4 <= 30\]\[2.50x + 14 <= 30 \]\[2.50x <= 30 - 14 \]\[2.50x <= 16 \]\[2.50 \div x \le 16\]

OpenStudy (secret-ninja):

Now you move the 2.50 to the other side and divide it into 16. \[x \le 16 \div 2.50\] \[x \le 6.4\]

OpenStudy (secret-ninja):

And you can't sell .4 of a burger, so he sold 6 hamburgers.

OpenStudy (secret-ninja):

I hope that this helped. @iam.dhue

OpenStudy (secret-ninja):

\(\bbox[8pt, pink, border:5pt solid green]{\Large Welcome~to~Openstudy!~:)}\) If this helped, please click the \(\large\sf\color{blue}{Best~Response}\) button and then click the \(\large\sf\color{blue}{Close}\) button to close your question so that you can post a new one later. Please feel free to tag me in any future questions you need help with by using the 'at' sign (@) and my user name. Like this: @Secret-Ninja \(\Huge\cal\color{darkgreen}{Have~a~great~day!}\)